# Graph-Structured Driven Dual Adaptation for Mitigating Popularity Bias

### v1

@@ V-A2 Baselines and Evaluation
To evaluate the effectiveness of  GSDA  in mitigating popularity bias, we implemented it using the state-of-the-art LightGCN framework [23]. We compared  GSDA  against several established debiasing baselines, including re-weighting models such as $\gamma$-AdjNorm [29] and DR-GNN [33], decorrelation approaches like MACR [10] and InvCF [34], as well as contrastive learning-based methods, including SGL [18], Adap-$\tau$ [35], and SimGCL [17]. Additionally, we benchmarked our method against the supervised alignment approach PAAC [19].
@@
To accurately evaluate the model’s ability to mitigate popularity bias, we adopted an unbiased dataset with uniformly distributed items in the test set, following established practices [37, 34, 28, 10]. This ensures that the evaluation focuses on the model’s effectiveness rather than its reliance on item popularity. Specifically, each item in the test set was assigned a fixed number of interactions, representing approximately 10% of the total dataset. To maintain the integrity of the test distribution, an additional 10% of interactions were randomly allocated to the validation set, while the remaining interactions were used for training. We evaluated Top-$K$ recommendation performance using $Recall@K$, $HR@K$, and $NDCG@K$. $Recall@K$ and $HR@K$ measure the model’s ability to provide comprehensive coverage, while $NDCG@K$ focuses on the quality of item ranking. To ensure fairness and avoid selection bias, we employed a full ranking strategy [38, 39], which evaluates all items rather than a subset [26]. Each experiment was repeated five times with different random seeds to account for variability, and the average scores were reported [26, 13]. This robust evaluation protocol ensures reliable and unbiased assessment of the model’s performance in mitigating popularity bias.
V-A3 Hyper-Parameter Settings.
@@ V-B Overall Performance
  • GSDA achieves substantial improvements compared to competitive baselines such as PAAC, Adap-$\tau$, and SimGCL, with performance gains ranging from 3.53% to 6.99%. For example, on the Movielens-10M dataset,  GSDA outperforms the best-performing baseline PAAC by 3.91% in NDCG@20 and by 5.30% in Recall@20. Similarly, on the Globo dataset,  GSDA achieves a 3.71% improvement in HR@20. Although PAAC demonstrates strong performance due to its supervised alignment strategy, which transfers information from popular items to unpopular ones,  GSDA surpasses PAAC by incorporating dual adaptive mechanisms. Specifically, our adaptive hierarchical alignment better captures distinct features at different layers, while our adaptive re-weighting contrastive learning dynamically adjusts sample weights, leading to superior representation learning.
@@
  • Our model  GSDA  is particularly effective in addressing data sparsity. For instance, on the Gowalla dataset, characterized by sparse user-item interactions,  GSDA achieves a notable 4.01% improvement in NDCG@20 and a 5.52% increase in Recall@20 over the strongest baseline. These improvements are attributed to our adaptive hierarchical alignment, which prioritizes shallow layers that retain distinct embedding features, and our adaptive re-weighting module, which dynamically balances contributions across items with varying levels of popularity. Collectively, these strategies ensure robust and discriminative embeddings, even in sparse data environments.
@@ V-C Ablation Study
  • GSDA-w/o SA: This variant entirely removes the adaptive hierarchical supervised alignment loss, thus eliminating the explicit alignment between popular and unpopular items. This setting isolates the contribution of the adaptive contrastive module, clarifying its independent effect on popularity bias mitigation.
  • GSDA-w/o CL: This variant omits the adaptive re-weighting contrastive loss, focusing solely on adaptive hierarchical supervised alignment. This setup allows us to independently assess how the hierarchical alignment contributes to enhancing item embedding distinctiveness.
  • GSDA-SAF: This variant replaces the adaptive layer-specific alignment strategy with a single fused alignment. Specifically, item embeddings aggregated from all GCN layers are aligned as a single unified representation rather than aligning embeddings layer-by-layer. This evaluates whether explicit layer-wise distinctions provide advantages over fused representations.
  • GSDA-SA0: This variant limits the supervised alignment process exclusively to the initial embedding layer ($\mathbf{h}_{i}^{0}$). By restricting alignment to the shallowest layer, we test the significance of incorporating richer structural information available in deeper layers.
  • GSDA-FRW: This variant employs fixed contrastive learning weights instead of dynamically computing them via the Gini coefficient. Static weights disregard the adaptive distribution of item popularity, allowing us to examine the necessity and impact of our adaptive weighting scheme.
@@
V-D Hyper-Parameter Sensitivities
In this section, we performed a comprehensive analysis of hyper-parameter sensitivities to validate and further understand the effectiveness of our proposed model. Specifically, we examined the influence of the alignment and contrastive learning coefficients ($\lambda_{1}$ and $\lambda_{2}$), the adaptive weighting parameter ($w$), and the impact of different grouping ratios for item popularity. Additionally, we analyzed the relationship between the number of hyperparameters and model performance, highlighting that our proposed adaptive strategies achieve superior recommendation accuracy with fewer hyperparameters compared to existing methods. Throughout our experiments, the dashed lines in each figure consistently denote the performance achieved by the strongest baseline. The results emphasize the importance of carefully balancing the adaptive hierarchical supervised alignment and adaptive re-weighting contrastive learning components, as well as the clear advantage of dynamically adjusting weights according to real-time item popularity distributions. Note that we primarily present experimental results on the Gowalla dataset due to its sparse interactions and representative nature. Our empirical findings remain consistent across other datasets.
V-D1 Effect of $\lambda_{1}$ and $\lambda_{2}$
V-D2 Effect of Adaptive Weighting ($w$)
V-D3 Effect of Group Ratio
V-D4 Effect of Hyperparameter Complexity
Figure 6: Comparison of model complexity and recommendation effectiveness. The horizontal axis denotes the number of hyperparameters required by each model, while the vertical axis represents their recommendation performance measured by NDCG@20 on the Gowalla dataset.
To further highlight the practical advantages of our dual-adaptive strategy, we compared  GSDA to the baselines in terms of both performance (NDCG@20) and the number of required hyperparameters, as depicted in Figure 6. Although PAAC achieves strong performance, it relies on a higher number of hyperparameters, increasing the complexity of model tuning and limiting practical usability. In contrast, our proposed  GSDA, benefiting from its adaptive hierarchical alignment and adaptive re-weighting mechanisms, achieves superior recommendation performance with significantly fewer hyperparameters. This not only simplifies hyperparameter tuning but also enhances the robustness and generalizability of our model across diverse datasets and scenarios, underscoring the practical benefits of adaptive parameterization strategies.
V-E Debias Ability.
@@ VI Related Work
GCN-based Recommender Systems. Graph Convolutional Networks (GCNs) have significantly advanced Collaborative Filtering (CF) by capturing high-order user-item interactions [24, 42]. Unlike traditional CF models, which primarily focus on first-order user-item relationships, GCN-based methods propagate and aggregate information through graph structures, enabling more expressive user and item representations.
@@
Despite these advancements, GCN-based methods still face challenges such as over-smoothing, where node representations become too similar as layers deepen, reducing model effectiveness [28, 29]. To address this, recent work explores residual connections [25], adaptive layer aggregation [23], and contrastive learning techniques [18] to enhance representation distinctiveness. However, balancing high-order propagation while maintaining feature diversity remains an open problem.
@@
Recent approaches address this by leveraging group-level alignment. Popularity-aware alignment [19] aligns items interacted with by the same users, allowing unpopular items to inherit characteristics from popular ones. Other techniques, such as DebiasCF [46], integrate supervised signals with contrastive learning to enhance representation robustness against bias. While these methods improve alignment, they often apply uniform constraints across all GCN layers, failing to consider the varying effects of deeper layers. Our approach introduces an adaptive alignment strategy that dynamically adjusts alignment strength per layer, improving representation learning while mitigating popularity bias [47, 48, 42].
Re-weighting for Mitigating Popularity Bias. Re-weighting strategies address popularity bias through static weighting, reinforcement learning, and contrastive learning. (1) Static Weighting Strategies assign fixed weights to items based on their popularity, aiming to counteract the tendency of recommendation models to favor popular items [49, 11, 50]. Methods like inverse propensity scoring (IPS) [51, 52] estimate item exposure probabilities and re-weight interactions accordingly, while MACR [10] employs counterfactual inference to down-weight frequently interacted items [53]. However, these approaches require careful calibration and struggle with dynamic shifts in popularity distributions. (2) Reinforcement Learning-based Approaches dynamically adjust item weights by treating the recommendation process as a decision-making problem [54, 32]. Neural Selection (NSE) [55] employs an actor-critic framework to optimize item selection, while DRSA [56] models weight assignment as a Markov Decision Process. These approaches improve adaptability but introduce high computational costs and instability in sparse data scenarios. (3) Contrastive Learning Variants aim to balance representation learning by treating popularity bias as a distributional issue [57, 58, 59, 60, 18]. Methods like Self-supervised Graph Learning (SGL) [18] augment training by randomly removing edges, while SimGCL [17] injects controlled noise to promote uniformity. However, such strategies may inadvertently amplify bias by enforcing uniform sampling across items.

### v2

@@ V-A2 Baselines and Evaluation
To evaluate the effectiveness of  GSDA  in mitigating popularity bias, we implemented it using the state-of-the-art LightGCN framework [23]. We compared  GSDA  against several established debiasing baselines, including re-weighting models such as $\gamma$-AdjNorm [29], DR-GNN [36] and LogDet [28], decorrelation approaches like MACR [10] and InvCF [37], as well as contrastive learning-based methods, including SGL [18], Adap-$\tau$ [38], and SimGCL [17]. Additionally, we benchmarked our method against the supervised alignment approach PAAC [19].
@@
To accurately evaluate the model’s ability to mitigate popularity bias, we adopted an unbiased evaluation protocol with a uniformly distributed item set in the test split, following established practices [40, 37, 28, 10]. In this setting, each item is assigned an equal number of interactions in the test set, thereby neutralizing the inherent advantage of frequently interacted (popular) items and preventing them from disproportionately influencing ranking metrics. This design ensures that any observed performance differences are attributable to the model’s capacity to handle items of varying popularity, rather than to skewed exposure frequencies. More concretely, each item in the test set is allocated a fixed number of interactions, representing approximately 10% of the total dataset. This constraint forces the model to perform well across the entire popularity spectrum, including long-tail items with historically low exposure, which is crucial for assessing bias mitigation effectiveness. To preserve the integrity of the evaluation distribution, an additional 10% of interactions are randomly assigned to the validation set, while the remaining interactions are used for training.
We evaluate Top-$K$ recommendation performance using $Recall@K$, $HR@K$, and $NDCG@K$. Here, $Recall@K$ and $HR@K$ reflect the ability to achieve broad coverage of relevant items, while $NDCG@K$ measures the ranking quality. Importantly, to avoid selection bias, we employ a full ranking evaluation strategy[41, 42]—ranking all candidate items rather than a sampled subset[26]—so that the results truly reflect the model’s ranking capability across the entire item set.
Finally, to ensure statistical reliability, each experiment is repeated five times with different random seeds, and average results are reported [26, 13]. This combination of an unbiased test distribution, full ranking evaluation, and repeated trials provides a robust and reproducible framework that faithfully measures a model’s ability to mitigate popularity bias without interference from pre-existing popularity effects.
V-A3 Hyper-Parameter Settings
@@ V-B Overall Performance
  • GSDA achieves substantial improvements compared to competitive baselines such as PAAC, Adap-$\tau$, and SimGCL, with performance gains ranging from 3.53% to 6.99%. For example, on the Movielens-10M dataset,  GSDA outperforms the best-performing baseline PAAC by 4.18% in NDCG@20 and by 5.47% in Recall@20. Similarly, on the Globo dataset,  GSDA achieves a 3.96% improvement in HR@20. Although PAAC demonstrates strong performance due to its supervised alignment strategy, which transfers information from popular items to unpopular ones,  GSDA surpasses PAAC by incorporating dual adaptive mechanisms. Specifically, our adaptive hierarchical alignment better captures distinct features at different layers, while our adaptive re-weighting contrastive learning dynamically adjusts sample weights, leading to superior representation learning.
@@
  • Our model  GSDA  is particularly effective in addressing data sparsity. For instance, on the Gowalla dataset, characterized by sparse user-item interactions,  GSDA achieves a notable 4.12% improvement in NDCG@20 and a 5.77% increase in Recall@20 over the strongest baseline. These improvements are attributed to our adaptive hierarchical alignment, which prioritizes shallow layers that retain distinct embedding features, and our adaptive re-weighting module, which dynamically balances contributions across items with varying levels of popularity. Collectively, these strategies ensure robust and discriminative embeddings, even in sparse data environments.
@@ V-C Ablation Study
  • GSDA-w/o SA: This variant entirely removes the adaptive hierarchical supervised alignment loss $L_{\mathrm{sa}}$, eliminating any explicit alignment between popular and unpopular items. It is designed to examine whether the adaptive contrastive module alone can effectively mitigate popularity bias.
  • GSDA-w/o CL: This variant omits the adaptive re-weighted contrastive loss $L_{\mathrm{cl}}$, focusing solely on adaptive hierarchical supervised alignment. It is intended to assess whether structural alignment alone can preserve semantic consistency between popular and unpopular items.
  • GSDA-SAF: This variant replaces the adaptive layer-specific alignment strategy with a single fused alignment, where item embeddings aggregated from all GCN layers are aligned as one unified representation. It is used to test whether treating all layers equally can match the effect of layer-wise alignment.
  • GSDA-SA0: This variant limits the supervised alignment process to the initial embedding layer ($\mathbf{h}_{i}^{0}$) only. It examines whether early alignment alone is sufficient without deeper-layer supervision.
  • GSDA-FRW: This variant uses fixed contrastive learning weights instead of dynamically computing them via the Gini coefficient. It is designed to evaluate the impact of ignoring popularity variations when setting contrastive weights.
@@
V-D Debias Ability
@@
V-E Hyper-Parameter Sensitivities
V-E1 Effect of $\lambda_{1}$ and $\lambda_{2}$
These results confirm that the two adaptive losses are complementary and need to be properly balanced to achieve optimal debiasing performance.
V-E2 Effect of Adaptive Weighting ($w$)
V-E3 Effect of Group Ratio
V-F Qualitative Analysis
Beyond quantitative results, we also provide a qualitative visualization of learned embeddings using t-SNE [45]. As shown in Figure 7, LightGCN embeddings tend to form dense clusters dominated by popular items, with many unpopular items collapsed into narrow regions. In contrast, GSDA achieves a more uniform dispersion of both popular and unpopular items, mitigating representation collapse and improving semantic separability. This observation provides intuitive support for our theoretical analysis and demonstrates the practical benefit of adaptive alignment in preserving representation diversity.
@@ VI Related Work
GCN-based Recommender Systems. Graph Convolutional Networks (GCNs) have significantly advanced Collaborative Filtering (CF) by capturing high-order user-item interactions [24, 46]. GCN-based methods propagate and aggregate information through graph structures, enabling more expressive user and item representations.
@@
Despite these advancements, GCN-based methods still face challenges such as over-smoothing, where node representations become too similar as layers deepen, reducing model effectiveness [28, 29]. To address this, recent work explores residual connections [25], adaptive layer aggregation [23], and contrastive learning techniques [18] to enhance representation distinctiveness. However, balancing high-order propagation while maintaining feature diversity remains an open problem. Beyond these architectural innovations, a line of theoretical and empirical studies further investigates why deeper GCNs tend to homogenize node representations and even lose expressivity as layers grow [31, 33, 32]. To alleviate such degradation, jump-connection style aggregation (e.g., JK-Net) has been proposed to adaptively select or combine information from multiple depths [48]. In recommender systems, recent analyses additionally reveal that the message-passing mechanism itself can amplify popularity bias with depth, disproportionately harming long-tail items [49, 50]. These findings collectively motivate the design of depth-aware architectures that preserve discriminability while leveraging high-order signals.
@@
Recent approaches address this by leveraging group-level alignment. Popularity-aware alignment [19] aligns items interacted with by the same users, allowing unpopular items to inherit characteristics from popular ones. Other techniques, such as DebiasCF [53], integrate supervised signals with contrastive learning to enhance representation robustness against bias. Complementary efforts regularize scoring or contextualize debiasing signals to reduce popularity-driven imbalance in predicted ranks [54, 55]. While these methods improve alignment, they often apply uniform constraints across all GCN layers, failing to consider the varying effects of deeper layers. At a broader level, fairness and stability perspectives in graph representation learning [56] highlight that popularity skew is a manifestation of distributional disparities, reinforcing the need for alignment strategies that remain stable across depths and user/item groups.In parallel, class-imbalance remedies such as the class-balanced loss and decoupled training are effective at re-balancing gradients and improving tail performance [57, 58]. Our approach introduces an adaptive alignment strategy that dynamically adjusts alignment strength per layer, improving representation learning while mitigating popularity bias [59, 60, 46].
Re-weighting for Mitigating Popularity Bias. Re-weighting strategies address popularity bias through static weighting, reinforcement learning, and contrastive learning. (1) Static Weighting Strategies assign fixed weights to items based on their popularity, aiming to counteract the tendency of recommendation models to favor popular items [61, 11, 62]. Methods like inverse propensity scoring (IPS) [63, 64] estimate item exposure probabilities and re-weight interactions accordingly, while MACR [10] employs counterfactual inference to down-weight frequently interacted items [65]. Foundational analyses have also documented the tension between accuracy and item-popularity effects [66], underscoring the need to explicitly manage long-tail distributions. However, these approaches require careful calibration and struggle with dynamic shifts in popularity distributions. (2) Reinforcement Learning-based Approaches dynamically adjust item weights by treating the recommendation process as a decision-making problem [67, 35]. Neural Selection (NSE) [68] employs an actor-critic framework to optimize item selection, while DRSA [69] models weight assignment as a Markov Decision Process. These approaches improve adaptability but introduce high computational costs and instability in sparse data scenarios. (3) Contrastive Learning Variants aim to balance representation learning by treating popularity bias as a distributional issue [70, 71, 72, 73, 18]. Methods like Self-supervised Graph Learning (SGL) [18] augment training by randomly removing edges, while SimGCL [17] injects controlled noise to promote uniformity. To reduce sampling and exposure bias in contrastive pipelines, debiased contrastive learning corrects the implicit negative sampling distribution [74]; more recently, recommendation-oriented debiased CL losses have been proposed to better cope with user–item sparsity and long-tail exposure [75]. However, such strategies may inadvertently amplify bias by enforcing uniform sampling across items.
@@ VII Conclusion
(a) LightGCN
(b) GSDA
Figure 7: t-SNE visualization of 400 randomly selected item embeddings.